# Astrometry meets Pulsar Timing Arrays: Synergies for Gravitational Wave Detection

### v1

@@ 3.2.2 A method for handling large Fisher matrices
In the examples discussed above we are able to perform the necessary computations involving Fisher matrices and their inversion, given the relatively small size of astrometric covariance matrix involved. However, for realistic datasets which may contain over $10^{6}$ stars, inverting matrices of such dimensionality can become demanding. To handle forecasts for such datasets, we develop a simple method to perform the Fisher forecasts in weak astrometric signal limit without needing to resort to the full numerical calculations. The weak signal limit is justified since current astrometric datasets are very much in the noise dominated regime and this is likely to be the case in the future as well. [^11] To be precise this requires $MI(f)\ll\sigma^{2}(f)$ where $\sigma^{2}$ denotes the frequency domain noise power spectral density. This condition ensures that the effect of the off-diagonal terms in the covariance is suppressed compared to the diagonal terms. Thus, we can approximate the matrix $D$ of equation (3.20) as
@@
$C_{\alpha}C^{-1}$ $=\begin{bmatrix}A_{\alpha}Q_{1}&\;\;\;\;B_{\alpha}Q_{2}\\ B^{T}_{\alpha}Q_{1}&\;\;\;\;D_{\alpha}Q_{2}\end{bmatrix}\begin{bmatrix}\mathds{1}_{N}&-B\sigma^{-2}_{S}\mathds{1}_{M}\\ -B^{T}A^{-1}&\mathds{1}_{M}\end{bmatrix}\,,$ (3.28)
$=\begin{bmatrix}A_{\alpha}Q_{1}-B_{\alpha}Q_{2}\,B^{T}A^{-1}&\;\;-\sigma_{S}^{-2}\,A_{\alpha}Q_{1}B+B_{\alpha}Q_{2}\\ B_{\alpha}^{T}Q_{1}-D_{\alpha}Q_{2}B^{T}A^{-1}&\;\;-\sigma_{S}^{-2}\,B_{\alpha}^{T}Q_{1}B+D_{\alpha}Q_{2}\end{bmatrix}\,.$
@@
[…]
@@
Figure 9: Fisher forecast for the amplitude and slope of the SGWB for a fiducial value $\log_{10}A_{\rm GW}=-14.6,\,\gamma=13/3$. In red, the PTA only case with the number of pulsars and noise parameters chosen to match the sensitivity of the combined IPTA dataset [9]. In blue, the forecasts on the addition of PTA–Astrometry cross-correlations with $10^{6}$ stars. Left: astrometry noise $\sigma=0.01$ mas. Right: astrometry noise $\sigma=0.02$ mas.
@@
PTA only $-14.6\pm 0.076$ $13/3\pm 0.22$
PTA + Astrometry, $\sigma=0.01$ mas $-14.6\pm 0.069$ $13/3\pm 0.19$
PTA + Astrometry, $\sigma=0.02$ mas $-14.6\pm 0.074$ $13/3\pm 0.21$
@@
We then use this approach to compute the full Fisher matrix, combining the individual per-frequency Fisher matrices. In Fig. 9, we present our results for a PTA setup with 75 pulsars and an astrometric setup with $10^{6}$ stars. The astrometric setup has $\Delta\theta_{\rm rms}=0.01\,\rm mas$ and $T_{\rm cad}=\mathrm{year}/15$, i.e. roughly corresponding to the observational cadence of Gaia. Our results indicate that for the case with $\sigma=0.01$ mas, we obtain roughly a $10\%$ improvement over the PTA-only constraints, while for $\sigma\geq 0.02$ mas, the improvement will be negligible. [^12] We note that for the $\sigma=0.01$ mas, the weak signal limit does not hold for the first frequency bin where the signal and noise are comparable in magnitude. However, our results from section 3.2.1 indicate that our perturbative expansion for the Fisher matrix may even underestimate the improvement compared to the full calculation (see fig. 8). See Table 1 for the exact numbers.
@@ 3.2.3 The case of dipolar anisotropy
$B_{pa}$ $=$ $\frac{(K_{p,a}^{(0)}+K_{p,a}^{(1)})I_{f}}{4\pi f}\,,$ (3.40)
$D_{ab}$ $=$ $(H_{ab}^{(0)}+H_{ab}^{(1)})I_{f}+\delta_{ab}\sigma^{2}_{a}\,.$ (3.41)
@@
PTA only $-14.6\pm 0.076$ $13/3\pm 0.22$ $0.05\pm 0.07$
PTA + Astrometry, $\sigma=0.01$ mas $-14.6\pm 0.069$ $13/3\pm 0.19$ $0.05\pm 0.029$
PTA + Astrometry, $\sigma=0.02$ mas $-14.6\pm 0.074$ $13/3\pm 0.21$ $0.05\pm 0.047$
@@
We choose the same fiducial $A_{\rm GW}$ and $\gamma$ as in section 3.2.1. We select a level of dipole anisotropy $\beta=5\times 10^{-2}$, finding this value to be the minimum level of kinematic dipole that will detectable with our PTA + Astrometry setup. Thus, astrometry can again help to tighten the constraints on the dipole anisotropy, which can be useful in determining the origin of the SGWB. Reaching the kinematic dipole level with PTA alone requires a substantial increase in the number of pulsars ($N>1000$[101, 88] which may only be achievable with SKA. It would be also interesting to explore how cross-correlations with astrometric datasets could work to improve SGWB sensitivity in the SKA-era.
@@ 4 Conclusions
In section 3 we studied the sensitivity of the astrometric and PTA setups to the SGWB, focusing on the measurement of the SGWB amplitude, spectral tilt and the magnitude of the kinematic dipole anisotropy. In section 3.1 we derived optimal estimators for SGWB monopole and dipole measurements in this context, built in terms of quadratic combinations of the astrometric deflections. We used these formulas to forecast the sensitivity of astrometry to the SGWB monopole and dipole, assuming a large number of stars uniformly distributed across the sky. In section 3.2, we analysed the joint PTA-Astrometry setup and showed that cross-correlating PTA and Astrometry data could tighten the constraints on SGWB obtained from PTA data alone. We used Fisher forecasts to calculate the sensitivity to the SGWB amplitude, spectral shape and dipole, finding that an astrometric survey with $0.01$ mas astrometric precision and the typical number of sources and cadence of Gaia could lead to noticeable improvements over current PTA only SGWB constraints. Improvements on the PTA constraints can be quite useful, since tighter constraints on the SGWB parameters can be used to rule out models and potentially distinguish between and astrophysical or cosmological signal.
@@ Acknowledgements
We are partially funded by the STFC grants ST/T000813/1 and ST/X000648/1. We also acknowledge the support of the Supercomputing Wales project, which is part-funded by the European Regional Development Fund (ERDF) via Welsh Government. We acknowledge the use of the GetDist software package for visualising the results of the Fisher forecasts [106]. For the purpose of open access, the authors have applied a Creative Commons Attribution licence to any Author Accepted Manuscript version arising.
@@ Appendix A Spectrum of angular deflection fluctuations
$\begin{split}\beta^{EB}&=\nabla_{i}\nabla_{b}^{\prime}\left[a_{1}\epsilon_{jab}n_{a}\left(\epsilon^{ilm}n_{l}q_{m}[{\bf(n\cdot q)}q^{j}-n^{j}]+\epsilon^{jlm}n_{l}q_{m}[{\bf(n\cdot q)}n^{i}-q^{i}]\right)\right]\\ &=\nabla_{i}\nabla_{b}^{\prime}\left[a_{1}T^{bi}\right]\,,\end{split}$ (A.3)
@@
$T^{bi}=$ $(q^{b}-({\bf n\cdot q})n^{b})(q^{i}+n^{i})(({\bf n\cdot q})-1)-(n^{i}q^{b}-({\bf n\cdot q})\delta^{ib})(({\bf n\cdot q})^{2}-1)\,,$ (A.4a)
and $a_{1}$ is defined in (2.17),
So we need $\nabla_{i}X,\nabla_{b}^{\prime}X,\nabla_{i}\nabla_{b}^{\prime}X$, where $X=a_{1},T^{bi}$. We can use the relations $\nabla_{i}n_{j}=\delta_{ij}-n_{i}n_{j}$, $\nabla_{i}^{\prime}q_{j}=\delta_{ij}-q_{i}q_{j}$, $\nabla_{i}q_{j}=\nabla_{i}^{\prime}n_{j}=\nabla_{i}v_{j}=\nabla_{i}^{\prime}v_{j}=0$, as well as
@@
$\begin{split}G(y)&=-\frac{2}{y^{2}(y-1)^{5}}\Bigg{[}(y-1)(-1+6y-42y^{2}+170y^{3}-481y^{4}+800y^{5}-684y^{6}+208y^{7}+96y^{8})\\ &\hskip 85.35826pt-3y^{2}(-1-y+13y^{2}-77y^{3}+198y^{4}-236y^{5}+128y^{6})\ln(y)\Bigg{]}\,,\end{split}$ (A.10)
@@ Appendix B Angular dependence of overlap reduction functions
$\left(-7\cos^{2}(\zeta)-2\cos(\zeta)+6(\cos(\zeta)-1)^{2}\ln\left(\sin^{2}\left(\frac{\zeta}{2}\right)\right)+5\right)^{2}$
@@
${\rm Tr}[{\boldsymbol{H}}_{1}{\boldsymbol{H}}_{1}]$ $=$ $\frac{64\pi^{2}\left(\alpha_{H}(Av)^{2}\sqrt{1-y}y^{2}+4\delta_{H}^{2}(y-1)y^{4}\left(\beta_{H}-\gamma_{H}\sqrt{1-y}\right)\right)}{9\left((nq)^{2}-1\right)^{2}(1-y)^{5/2}}$ (B.3)
@@
$6(y-1)y^{2}\left(2\left(nq^{2}+1\right)y^{2}+nq(nq+8)y-2nq+y\right)\ln(y)-9(nq+1)^{2}y^{4}\ln^{2}(y)$
@@
$-(nv)\sqrt{1-y}\sqrt{4\left((nv)^{2}-1\right)(y-1)y-(Av)^{2}}$
$\gamma_{H}$ $=$ $(nv)^{2}\left((nq)^{2}-2y+1\right)+(nq)^{2}(-y)+y$ (B.7)
@@
$\left(\cos^{2}(\zeta)-\cos(\zeta)-\ln(8)\cos(\zeta)+3(\cos(\zeta)-1)\ln(1-\cos(\zeta))-2+\ln(8)\right)^{2}$

### v2

@@ 3.2.2 A method for handling large Fisher matrices
In the examples discussed above we are able to perform the necessary computations involving Fisher matrices and their inversion, given the relatively small size of astrometric covariance matrix involved. However, for realistic datasets which may contain over $10^{6}$ stars, inverting matrices of such dimensionality can become demanding. To handle forecasts for such datasets, we develop a simple method to perform the Fisher forecasts in weak astrometric signal limit without needing to resort to the full numerical calculations. The weak signal limit is justified since current astrometric datasets are very much in the noise dominated regime and this is likely to be the case in the future as well. [^17] To be precise this requires $MI(f)\ll{4\pi}\sigma^{2}(f)$ where $\sigma^{2}$ denotes the frequency domain noise power spectral density. This condition ensures that the effect of the off-diagonal terms in the covariance is suppressed compared to the diagonal terms. Thus, we can approximate the matrix $D$ of equation (3.27) as
@@
$C_{\alpha}C^{-1}$ $=\begin{bmatrix}A_{\alpha}Q_{1}&\;\;\;\;B_{\alpha}Q_{2}\\ B^{T}_{\alpha}Q_{1}&\;\;\;\;D_{\alpha}Q_{2}\end{bmatrix}\begin{bmatrix}\mathds{1}_{N}&-B\sigma^{-2}_{S}\mathds{1}_{M}\\ -B^{T}A^{-1}&\mathds{1}_{M}\end{bmatrix}\,$ (3.34)
$=\begin{bmatrix}A_{\alpha}Q_{1}-B_{\alpha}Q_{2}\,B^{T}A^{-1}&\;\;-\sigma_{S}^{-2}\,A_{\alpha}Q_{1}B+B_{\alpha}Q_{2}\\ B_{\alpha}^{T}Q_{1}-D_{\alpha}Q_{2}B^{T}A^{-1}&\;\;-\sigma_{S}^{-2}\,B_{\alpha}^{T}Q_{1}B+D_{\alpha}Q_{2}\end{bmatrix}\,$,
@@
${\cal M}_{\alpha}^{(0)}$ $=$ $\begin{bmatrix}A_{\alpha}\,A^{-1}&\,\,0\\ B_{\alpha}^{T}\,A^{-1}&\,\,0\end{bmatrix}\,$, (3.41)
@@
Figure 8: Fisher forecast for the amplitude and slope of the SGWB for a fiducial value $\log_{10}A_{\rm GW}=-14.6,\,\gamma=13/3$. In red, the PTA only case with the number of pulsars and noise parameters chosen to match the sensitivity of the combined IPTA dataset [9]. In blue, the forecasts on the addition of PTA–Astrometry cross-correlations with $10^{6}$ stars. Left: astrometry noise $\sigma=0.001$ mas. Right: astrometry noise $\sigma=0.01$ mas.
@@
PTA only $14.6\pm 0.079$ $13/3\pm 0.24$
PTA + Astrometry, $\sigma=0.001$ mas $14.6\pm 0.072$ $13/3\pm 0.19$
PTA + Astrometry, $\sigma=0.01$ mas $14.6\pm 0.079$ $13/3\pm 0.24$
@@
We now use this approach to compute the full Fisher matrix, combining the individual per-frequency Fisher matrices. In Fig. 8, we present our results for a PTA setup with 75 pulsars and an astrometric setup with $10^{6}$ stars. The astrometric setup has $\Delta\theta_{\rm rms}=0.01,\,0.001\,\rm mas$ and $T_{\rm cad}=\mathrm{year}/15$, i.e. roughly corresponding to the observational cadence of Gaia. Our results indicate that for the case with $\sigma=0.001$ mas, we obtain roughly a $\mathcal{O}(10\%)$ improvement over the PTA-only constraints, while for $\sigma\geq 0.01$ mas, the improvement will be negligible. [^18] We note that for the $\sigma=0.001$ mas case, the weak signal limit does not hold for the first few frequency bins where the signal and noise are comparable in magnitude. However, we have checked (with lower number of stars) that for similar noise vs signal strengths, the results from the series expansion match fairly well with the full numerical results. See Table 1 for the exact numbers.
@@ 3.2.3 The case of dipolar anisotropy
$B_{pa}$ $=$ $\frac{(K_{p,a}^{(0)}+K_{p,a}^{(1)})I_{f}}{{(4\pi)^{2}}f}\,$, (3.46)
$D_{ab}$ $=$ $\frac{(H_{ab}^{(0)}+H_{ab}^{(1)})}{{4\pi}}I_{f}+\delta_{ab}\sigma^{2}_{a}\,.$ (3.47)
@@
PTA only $14.6\pm 0.079$ $13/3\pm 0.24$ $0.05\pm 0.081$
PTA + Astrometry, $\sigma=0.001$ mas $14.6\pm 0.072$ $13/3\pm 0.18$ $0.05\pm 0.048$
PTA + Astrometry, $\sigma=0.01$ mas $14.6\pm 0.079$ $13/3\pm 0.24$ $0.05\pm 0.080$
@@
We choose the same fiducial $A_{\rm GW}$ and $\gamma$ as in section 3.2.1. We select a level of dipole anisotropy $\beta=5\times 10^{-2}$, finding this value to be the minimum level of kinematic dipole that will detectable with our PTA + Astrometry setup. Thus, astrometry can again help to tighten the constraints on the dipole anisotropy, which can be useful in determining the origin of the SGWB. Given that the sensitivity to the SGWB is likely to be dominated by PTA datasets even in the near future, one would likely require futuristic PTA experiments combined with futuristic astrometric surveys to reach the CMB kinematic dipole level. Reaching the CMB kinematic dipole level ($\beta=1.23\times 10^{-3}$) with PTA alone requires a substantial increase in the number of pulsars (N>1000N>1000[102, 88] which may only be achievable with SKA. It would be also interesting to explore how cross-correlations with astrometric datasets could work to improve SGWB sensitivity in the SKA-era.
@@ 4 Conclusions
In section 3 we studied the sensitivity of the astrometric and PTA setups to the SGWB, focusing on the measurement of the SGWB amplitude, spectral tilt and the magnitude of the kinematic dipole anisotropy. In appendix C we derived optimal estimators for SGWB monopole and dipole measurements in this context, built in terms of quadratic combinations of the astrometric deflections. We used these formulas to forecast the sensitivity of astrometry to the SGWB monopole and dipole, assuming a large number of stars uniformly distributed across the sky. In section 3.2, we analysed the joint PTA-Astrometry setup and showed that cross-correlating PTA and Astrometry data could tighten the constraints on SGWB obtained from PTA data alone. We used Fisher forecasts to calculate the sensitivity to the SGWB amplitude, spectral shape and dipole, finding that an astrometric survey with $0.01$ mas astrometric precision and the typical number of sources and cadence of Gaia could lead to noticeable improvements over current PTA only SGWB constraints. Improvements on the PTA constraints can be quite useful, since tighter constraints on the SGWB parameters can be used to rule out models and potentially distinguish between and astrophysical or cosmological signal. Reaching the level of sensitivity required to detect the CMB level kinematic dipole anisotropy in the SGWB is likely to require both futuristic PTA experiments such as SKA, as well as futurustic astrometric surveys.
@@ Acknowledgements
We are partially funded by the STFC grants ST/T000813/1 and ST/X000648/1. We also acknowledge the support of the Supercomputing Wales project, which is part-funded by the European Regional Development Fund (ERDF) via Welsh Government. We acknowledge the use of the GetDist software package for visualising the results of the Fisher forecasts [107]. The codes used to perform the numerical calculations and plot the results in this paper are publicly available on GitHub. For the purpose of open access, the authors have applied a Creative Commons Attribution licence to any Author Accepted Manuscript version arising.
@@ Appendix A Spectrum of angular deflection fluctuations
[…]
@@
$T^{bi}=(q^{b}-({\bf n\cdot q})n^{b})(q^{i}+n^{i})(({\bf n\cdot q})-1)-(n^{i}q^{b}-({\bf n\cdot q})\delta^{ib})(({\bf n\cdot q})^{2}-1)\,$, (A.4)
and $a_{1}$ is defined in (2.19). So we need $\nabla_{i}X,\nabla_{b}^{\prime}X,\nabla_{i}\nabla_{b}^{\prime}X$, where $X=a_{1},T^{bi}$. We can use the relations $\nabla_{i}n_{j}=\delta_{ij}-n_{i}n_{j}$, $\nabla_{i}^{\prime}q_{j}=\delta_{ij}-q_{i}q_{j}$, $\nabla_{i}q_{j}=\nabla_{i}^{\prime}n_{j}=\nabla_{i}v_{j}=\nabla_{i}^{\prime}v_{j}=0$, as well as
@@
[…]
@@ Appendix B Angular dependence of overlap reduction functions
$\left(-7\cos^{2}(\zeta)-2\cos(\zeta)+6(\cos(\zeta)-1)^{2}\ln\left(\sin^{2}\left(\frac{\zeta}{2}\right)\right)+5\right)^{2}\,;$
@@
${\rm Tr}[{\boldsymbol{H}}_{1}{\boldsymbol{H}}_{1}]$ $=$ $\frac{64\pi^{2}\left(\alpha_{H}(Av)^{2}\sqrt{1-y}y^{2}+4\delta_{H}^{2}(y-1)y^{4}\left(\beta_{H}-\gamma_{H}\sqrt{1-y}\right)\right)}{9\left((nq)^{2}-1\right)^{2}(1-y)^{5/2}}\,;$ (B.3)
@@
$6(y-1)y^{2}\left(2\left(nq^{2}+1\right)y^{2}+nq(nq+8)y-2nq+y\right)\ln(y)-9(nq+1)^{2}y^{4}\ln^{2}(y)\,$,
@@
$-(nv)\sqrt{1-y}\sqrt{4\left((nv)^{2}-1\right)(y-1)y-(Av)^{2}}\,$,
$\gamma_{H}$ $=$ $(nv)^{2}\left((nq)^{2}-2y+1\right)+(nq)^{2}(-y)+y\,$, (B.7)
@@
$\left(\cos^{2}(\zeta)-\cos(\zeta)-\ln(8)\cos(\zeta)+3(\cos(\zeta)-1)\ln(1-\cos(\zeta))-2+\ln(8)\right)^{2}\,$,
@@
Appendix C Optimal Estimators
In this section we briefly discuss how to build optimal estimators for the coefficients $p_{n}$ formally appearing in the sum (3.1), later specialising to the monopole and dipole case with $p_{0}$ and $p_{1}$. The determination of these coefficients allows us to infer the properties of the SGWB. First, we recall here that we can expand the astrometric deflection correlation in terms of these basis coefficients $p_{n}$ as
$\langle{\delta n_{a}^{i}\,\delta n_{b}^{j}}\rangle\equiv C^{ij}_{ab}=\sum_{n}p_{n}H_{ab,n}^{ij}+N_{ab}^{ij}\,$, (C.1)
in terms of the response $H$ and the noise $N$. Now, following the approach previously developed for the CMB [108], a quadratic estimator for $p_{n}$ can be formally expressed as
$\hat{p}_{n}=\delta n^{i}_{a}E_{ab,n}^{ij}\delta n^{j}_{b}-b_{n}\,$, (C.2)
in terms of a symmetric matrix $E_{ab,n}$ and a vector $b_{n}$.
We wish to determine the corresponding values of these quantities which render the estimator unbiased i.e. $\langle{\hat{p}_{n}}\rangle=p_{n}$, and minimize its variance. Taking the expectation value of the estimator we find
$\langle{\hat{p}_{n}}\rangle=\sum_{n^{\prime}}W_{nn^{\prime}}p_{n^{\prime}}+{\rm Tr}[{\boldsymbol{E}}_{n}{\boldsymbol{N}}]-b_{n},\qquad{\rm with}\qquad W_{nn^{\prime}}\equiv{\rm Tr}[{\boldsymbol{H}}_{n^{\prime}}{\boldsymbol{E}}_{n}]\,.$ (C.3)
Thus, the quantity $b_{n}$ should be chosen as $b_{n}={\rm Tr}[{\boldsymbol{E}}_{n}{\boldsymbol{N}}]$ to ensure that the estimator is unbiased. The covariance of matrix of the estimators $\mathrm{Cov}(\hat{p}_{n},\hat{p}_{n^{\prime}})\equiv\langle{(\hat{p}_{n}-\langle{\hat{p}_{n}}\rangle)(\hat{p}_{n^{\prime}}-\langle{\hat{p}_{n^{\prime}})}\rangle}\rangle$, assuming that the astrometric deflections obey a Gaussian distribution, can be written as
$\mathrm{Cov}(\hat{p}_{n},\hat{p}_{n^{\prime}})$ $=\left[C_{ad}^{il}C_{bc}^{jk}+C_{ac}^{ik}C_{bd}^{kl}\right]E^{ij}_{ab,n}E^{kl}_{cd,n^{\prime}}$ (C.4)
$=2{\rm Tr}[{\boldsymbol{C}}{\boldsymbol{E}}_{n}{\boldsymbol{C}}{\boldsymbol{E}}_{n^{\prime}}]\,.$ (C.5)
The solution for the matrix ${\boldsymbol{E}}_{n}$ which minimises this variance under the normalisation constraint $W_{nn}=1$ is given by [108]
${\boldsymbol{E}}_{n}=\frac{{\boldsymbol{C}}^{-1}{\boldsymbol{H}}_{n}{\boldsymbol{C}}^{-1}}{{\rm Tr}[{\boldsymbol{C}}^{-1}{\boldsymbol{H}}_{n}{\boldsymbol{C}}^{-1}{\boldsymbol{H}}_{n}]}\,$, (C.6)
so that
$W_{nn^{\prime}}=\frac{{\rm Tr}\left[{\boldsymbol{C}}^{-1}{\boldsymbol{H}}_{n^{\prime}}{\boldsymbol{C}}^{-1}{\boldsymbol{H}}_{n}\right]}{{\rm Tr}\left[{\boldsymbol{C}}^{-1}{\boldsymbol{H}}_{n}{\boldsymbol{C}}^{-1}{\boldsymbol{H}}_{n}\right]}\,.$ (C.7)
The variance of the individual $\hat{p}_{n}$ is thus
${\text{Var}}(\hat{p}_{n})=\frac{2}{{\rm Tr}\left[{\boldsymbol{C}}^{-1}{\boldsymbol{H}}_{n}{\boldsymbol{C}}^{-1}{\boldsymbol{H}}_{n}\right]}\,.$ (C.8)
This estimator is optimal in the sense that it is unbiased and that its variance is equal to the what we obtained using the Fisher matrix approach in section 3.1, making it the minimal variance estimator that saturates the Cramer-Rao bound (see [99, 108] for a detailed discussion).
The expressions obtained here are completely general and valid for arbitrary distributions of stars with arbitrary noise properties which are encoded in the noise matrix $\boldsymbol{N}$. Further simplifications can be made in the noise dominated limit with the diagonal noise where the solution reduces to
${\boldsymbol{E}}_{n}=\frac{{\boldsymbol{H}}_{n}}{{\rm Tr}[{\boldsymbol{H}}_{n}{\boldsymbol{H}}_{n}]},\quad W_{nn^{\prime}}=\frac{{\rm Tr}[{\boldsymbol{H}}_{n}{\boldsymbol{H}}_{n^{\prime}}]}{{\rm Tr}[{\boldsymbol{H}}_{n}{\boldsymbol{H}}_{n}]}\,.$ (C.9)
Finally, in the special case where one only wishes to estimate a single $p_{n}$, e.g. the monopole, from the data, the estimators become trivial
$\hat{p}_{n}=\frac{\delta n^{i}_{a}H_{ab,n}^{ij}\delta n_{b}^{j}}{({\rm Tr}[{\boldsymbol{H}}_{n}{\boldsymbol{H}}_{n}])}\,-b_{n}.$ (C.10)